fee? 5) Amon g coffee drinkers, men drink a mean of 3.4 cups per day with a standard deviation of 0.8 cup. Assum distribution. e the number of cups per day follows a normal a. What proportion drinks 2 cups per day or more? b. What proportion drinks no more than 4 cups per day?
media%2Fb80%2Fb80fc955-aff1-40fe-b008-2f
Answer #1

5a)

for normal distribution z score =(X-μ)/σ
here mean=       μ= 3.2
std deviation   =σ= 0.800
probability = P(X>2) = P(Z>-1.5)= 1-P(Z<-1.5)= 1-0.0668= 0.9332

b)

probability = P(X>4) = P(Z>1)= 1-P(Z<1)= 1-0.8413= 0.1587

c)

for top 5% critical z=1.645

hence corresponding number of coffee cups=mean+z*std deviation=3.2+1.645*0.8 =4.52

d)

sample size       =n= 20
std error=σ=σ/√n= 0.1789
probability = P(X>3) = P(Z>-1.12)= 1-P(Z<-1.12)= 1-0.1314= 0.8686